# appendix,by

### v1

@@ 2.1 Purity
    The formulas saying that $(\operatorname{res}(y_{jh}))_{j\in[1,n],h\in[r,s+\ell]}$ is a $k$-generic point of the irreducible algebraic variety $\operatorname{in}_{0}(H_{r,s})$.
Let $\Delta$ be a finite fragment of $\Theta$. Without loss of generality, $\Delta$ says that a finite subtuple of variables $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ lies in a certain Zariski open dense subset $H^{\prime}$ of $H_{r,s}$, that all of its coordinates have valuation 0, and that its residue lies in a certain Zariski open dense subset $H^{\prime\prime}$ of $\operatorname{in}_{0}(H_{r,s})$. Since the residue field $k$ is algebraically closed, there is a point $(a_{jh})_{j\in[1,n],h\in[r,s+\ell]}\in H^{\prime\prime}(k)$, and by Proposition˜1.13 the set of points of valuation 0 in $H_{r,s}(K)$ with residue $(a_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ is Zariski dense in $H_{r,s}$, so in particular there is such a point $(\alpha_{jh})_{j\in[1,n],h\in[r,s+\ell]}\in H^{\prime}(K)$. Therefore $\Theta$ is finitely satisfiable, so by the Compactness Theorem it is consistent.
@@
Therefore, extending the action of $\sigma_{\to}$ from polynomials to formulas in the obvious way, as $(\alpha_{jh})_{j\in[1,n],h\in\mathbb{Z}}$ realises $\Theta$, for every $r\leq s\in\mathbb{Z}$ we have, in $\mathcal{L}_{\mathrm{vf}}$,
@@
Recall that, in an $R$-module, a partial positive primitive-type (or partial pp-type) is specified by a filter of pp-definable cosets, hence can be written by using at most $\lvert R\rvert+\aleph_{0}$-many formulas (see e.g. [Pre88] for model theory of modules). In particular, under our assumptions $\mathcal{O}^{\times}$ realises all consistent partial pp-types. By ˜2.14 the short exact sequence of $\mathbb{Z}[\sigma]$-modules $1\to\mathcal{O}^{\times}\to K^{\times}\to\Gamma\to 0$ is pure, hence the existence of a cross-section follows from [Pre88, Theorem 2.8]. As standard, a cross-section $s$ induces an angular component $\operatorname{ac}_{s}(x)\coloneqq\operatorname{res}(x/s(v(x)))$, and clearly if $s$ commutes with $\sigma$ then so does $\operatorname{ac}_{s}$, since both $\operatorname{res}$ and $v$ do. The “in particular” part follows from ˜2.2. ∎
@@ 2.2 Good reducts
Let $T_{0}$ be an inductive theory that is not model complete (e.g. the theory of oags with an automorphism, or the theory of fields), and let $T_{1}$ be its Morleyisation (see the references given at the start of Subsection˜4.3). Observe that every model of $T_{0}$ expands to a model of $T_{1}$, and that every model of $T_{1}$ is existentially closed by quantifier elimination. In particular, if $M_{0}\vDash T_{0}$ is not existentially closed, then its natural expansion to $M_{1}\vDash T_{1}$ is existentially closed but has a non-existentially closed reduct.
@@
Let $T$ be the $\mathbb{Z}[\sigma]$-submodule of $L^{\times}$ generated by $K^{\times}$ and $\mathcal{O}(L)^{\times}$, and let $\theta\colon T\rightarrow k(L)^{\times}$ be the unique morphism which coincides with $\operatorname{ac}$ on $K^{\times}$ and with $\operatorname{res}$ on $\mathcal{O}(L)^{\times}$. Enumerate $T$ in a tuple $\bar{b}$. If we prove that the pp-type of $\bar{b}$ in the $\mathbb{Z}[\sigma]$-module $L^{\times}$ is contained in the pp-type of $\theta(\bar{b})$ in $k(L)^{\times}$, then we may conclude by using existential saturation of $k(L)^{\times}$ as a $\mathbb{Z}[\sigma]$-module to apply Point (v) in [Pre88, Theorem 2.8], obtaining a morphism $\operatorname{ac}_{L}\colon L^{\times}\rightarrow k(L)^{\times}$ which extends $\theta$.
@@
For every $r\leq s\in\mathbb{Z}$, there is a group homomorphism $\psi_{r,s}\colon\mathbb{G}_{\mathrm{m}}^{n(s+\ell-r+1)}\rightarrow\mathbb{G}_{\mathrm{m}}^{d(s-r+1)}$ satisfying $\psi_{r,s}(\operatorname{orb}_{r,s+\ell}(\alpha))=\operatorname{orb}_{r,s}(a)$, namely we can take the homomorphism which maps $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ to
@@
Let $e=(e_{1},\ldots,e_{d})$ denote the tuple $\xi_{0,0}(\operatorname{orb}_{0,\ell}(\alpha))$. Consider the set of $\mathcal{L}_{\mathrm{field}}$-formulas $\Omega$ over $k(L)$ in variables $(y_{jh})_{j\in[1,n],h\in\mathbb{Z}}$ which says that each finite subtuple $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ is a $k(L)$-generic point of the algebraic variety $V_{r,s}$ defined by the equations $\xi_{r,s}((z_{jh})_{j\in[1,n],h\in[r,s+\ell]})=\theta(\operatorname{orb}_{r,s}(e))$.
To see that $\Omega$ is consistent, consider a finite fragment mentioning formulas in the finite subtuple of variables $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$. As $k(L)^{\times}$ is a divisible abelian group, there is a (possibly non-equivariant) angular component $\operatorname{ac}^{\prime}\colon L\to k(L)$ extending $\theta$. Since, for every $r\leq s$, we have $\operatorname{orb}_{r,s}(e)\in\operatorname{im}(\xi_{r,s})$, and $\operatorname{ac}^{\prime}$ is a multiplicative homomorphism, $V_{r,s}(k(L))$ is not empty. As $V_{r,s}(k(L))$ is a coset of $\operatorname{in}_{0}(\ker\xi_{r,s})$ (which does not depend on the choice of an angular component by Lemma˜1.14), by Lemma˜2.11 it is irreducible, and it follows that $\Omega$ is consistent.
Hence the set $\Omega$ is realised in some field extension of $k(L)$; let $(\beta_{jh})_{j\in[1,n],h\in\mathbb{Z}}$ be a realisation, and consider the field generated $k(L)((\beta_{jh})_{j\in[1,n],h\in\mathbb{Z}})$. We now have that for every $r\leq s$, the finite tuple $(\beta_{j(h+1)})_{j\in[1,n],h\in[r,s+\ell]}$ is generic in the algebraic variety defined by the equations
@@
As $\mathsf{ACVF}_{\mathrm{ac}}$ eliminates quantifiers [^9] This is folklore, and follows from e.g. the much more general [HH19, Theorem 1]. For a direct proof, use [HKR18, Fact 2.4] in combination with [Tou23, Observation 1.51]., arguing as in the proof of Proposition˜2.3(1) we see that $K$ is algebraically closed, and in particular closed under taking roots. Suppose that $(L_{0},v_{L_{0}},\sigma_{L_{0}})\supseteq(K,v,\sigma)$ is such that $L_{0}\vDash\exists x\;\varphi(x)$, with $\varphi(x)$ a quantifier-free $\mathcal{L}_{\sigma}(K)$-formula. Let $(L,v_{L},\sigma_{L})$ be an existentially $\aleph_{1}$-saturated existentially closed extension of $L_{0}$. After applying Proposition˜2.18, $L$ extends $K$ as $\mathrm{ac}$-valued difference fields, and $L\vDash\exists x\;\varphi(x)$. Since $\mathcal{L}_{\mathrm{ac},\sigma}\supseteq\mathcal{L}_{\sigma}$ and $(K,v,\sigma,\operatorname{ac})$ is existentially closed, we have $K\vDash\exists x\;\varphi(x)$. ∎
@@
It is well known that the theory $\mathsf{ACVF}_{s}$ of algebraically closed valued fields with a cross-section eliminates quantifiers; see e.g. [Kes24, Theorem 2.2] for a proof for $\mathsf{ACVF}_{s,\iota}$ that can be easily adapted.
@@
Consider the $\mathbb{Z}[\sigma]$-submodule $N$ of $L$ consisting of elements of $L$ with angular component 1; note that the image of $s$ is contained in $N$. Let $\bar{b}$ be a tuple enumerating $\Gamma(K)$. If we prove that the pp-type of $\bar{b}$ in $\Gamma(L)$ is contained in the pp-type of $s(\bar{b})$ in $N$, then we may apply [Pre88, Theorem 2.8(v)] and obtain the desired extension of $s$ to $\Gamma(L)$. In other words, we need to show that if $a$ is a finite subtuple of $\bar{b}$, and $A\in\operatorname{Mat}_{d\times n}(\mathbb{Z}[\sigma])$ is such that $A\cdot z=a$ has a solution in $\Gamma(L)$, then $A\cdot z=s(a)$ has a solution in $N$. In fact, it suffices to show that $A\cdot z=a$ has a solution in $\Gamma(K)$, as then if $\gamma$ is such a solution we get that $A\cdot s(\gamma)=s(a)$ and $s(\gamma)\in N$.
This can follows from Lemma˜2.20, concluding the proof. ∎
@@
Let $K$ be an existentially closed $\operatorname{ac}$-valued (resp., $s$-valued) difference field. Then $\operatorname{VF}(K)$ is algebraically closed and $(\Gamma,\sigma)$ and $(k,\sigma)$ are existentially closed in the respective categories. In particular, $(K,v,\operatorname{ac})$ (resp., $(K,v,s)$) and $(K,v)$ are existentially closed.
@@
By Proposition˜2.21 it suffices to deal with the $\operatorname{ac}$-valued case. By Corollary˜2.19 $(K,v,\sigma)$ is existentially closed, so we may apply Proposition˜2.3. Alternatively, we may adapt the proof of the latter by using quantifier elimination in $\mathsf{ACVF}_{\mathrm{ac}}$. The “in particular” part follows from the above and Proposition˜2.3(1). ∎
@@ 2.3 Bad reducts
Another natural reduct of $(K,v,\sigma)$ to consider is the difference field $(K,\sigma)$. In this subsection we will see that, contrary to what we saw in Subsection˜2.2, forgetting the valuation does not preserve existential closedness.
Recall that existentially closed difference fields are axiomatisable by a theory denoted by $\mathsf{ACFA}$, in the language $\{+,0,-,\cdot,1,(-)^{-1},\sigma,\sigma^{-1}\}$. This theory can be axiomatised by saying that $(K,\sigma)\vDash\mathsf{ACFA}$ iff $K$ is algebraically closed and, for all $n$, all irreducible algebraic subvarieties $V,W$ of $K^{n}$ and $K^{2n}$ respectively such that $W\subseteq V\times\sigma(V)$ and the projections of $W$ to $V$ and to $\sigma(V)$ are dominant, there is $x\in V(K)$ such that $(x,\sigma(x))\in W$. See [Mac97, Theorem 7], [CH99, Theorem on p. 3007].
We will see in ˜3.30 that if $k\vDash\mathsf{ACFA}$ has characteristic 0 and $\Gamma$ is an existentially closed difference oag, then the Hahn field $k((t^{\Gamma}))$, made into a valued difference field as in ˜1.6, is existentially closed. The following argument, due to Philip Dittmann, shows that it is not existentially closed as a difference field, that is, it is not a model of $\mathsf{ACFA}$.
@@
It is well known (see e.g. [CH99, Proposition 1.2]) that the fixed field of every model of $\mathsf{ACFA}$ is pseudofinite, and in particular not separably closed. By [FJ23, Corollary 12.5.5] [^10] In the third edition of the same book (2008), this is Corollary 11.5.5. every pseudo-algebraically closed (in particular, every pseudofinite) nontrivially valued henselian field is separably closed. Since there is a fixed point of nonzero valuation, the induced valuation on the fixed field $K_{0}$ of $K$ is nontrivial. Hence, to obtain a contradiction, it suffices to show that $K_{0}$ is henselian.
We check henselianity directly. Let $f\in\mathcal{O}(K_{0})[x]$ and $a\in\mathcal{O}(K_{0})$ be such that $\operatorname{res}(a)$ is a simple root of $\operatorname{res}(f)$. By henselianity of $K$ there is $b\in\mathcal{O}(K)$ with $f(b)=0$ and v(ba)>0v(b-a)>0. If $\sigma(b)=b$ we are done, so assume this is not the case. As $f$ is over the fixed field, $\sigma(b)$ is a root of $f$. Therefore $(x-b)(x-\sigma(b))$ divides $f$, and since $\operatorname{res}(\sigma(b))=\sigma(\operatorname{res}(b))=\sigma(\operatorname{res}(a))=\operatorname{res}(a)$ this contradicts that the latter is a simple root of $\operatorname{res}(f)$. ∎
@@
For notational simplicity, we will only deal with the case $n=2$. Define $P(x)\coloneqq(x-a)(x-\sigma(a))$, observe that it is over the fixed field $K_{0}$, and set $Q(x,y)\coloneqq P(x)P(y)-\varepsilon$. By the Eisenstein criterion (in the version of [FJ23, Lemma 2.3.10(b)]) applied to $(K[x])[y]$, the vanishing variety $V$ of $Q$ is absolutely irreducible.
@@
Let us show that every existentially closed valued difference field satisfies the assumptions of Proposition˜2.26. The existence of $\varepsilon$ follows from existential closedness of $\Gamma$ (Proposition˜2.3(2)). To find $a$ with $v(a)=0$, $\sigma^{2}(a)=a$, and $\operatorname{res}(\sigma(a))\neq\operatorname{res}(a)$, it suffices to take two Gauss extensions (see [EP05, Corollary 2.2.2]), let $\sigma$ swap the added points, and apply existential closedness. ∎
@@
The proof above may be adapted, for instance, to the existentially closed isometric case studied in [AD11], but not to the existentially closed strongly contractive case, where the valued field sort is indeed a model of $\mathsf{ACFA}$ [CH14, Fact 2.13].

### v2

@@ 2.1 Purity
    The formulas saying that $(\operatorname{res}(y_{jh}))_{j\in[1,n],h\in[r,s+\ell]}$ is a $k$-generic point of the irreducible variety $\operatorname{in}_{0}(H_{r,s})$.
Let $\Delta$ be a finite fragment of $\Theta$. Without loss of generality, $\Delta$ says that a finite subtuple of variables $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ lies in a certain Zariski open dense subset $H^{\prime}$ of $H_{r,s}$, that all of its coordinates have valuation 0, and that its residue lies in a certain Zariski open dense subset $H^{\prime\prime}$ of $\operatorname{in}_{0}(H_{r,s})$. Since the residue field $k$ is algebraically closed, there is a point $(a_{jh})_{j\in[1,n],h\in[r,s+\ell]}\in H^{\prime\prime}(k)$, and by Proposition˜1.13 the set of points of valuation 0 in $H_{r,s}(K)$ with residue $(a_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ is Zariski dense in $H_{r,s}$, so in particular there is such a point in $H^{\prime}(K)$. Therefore $\Theta$ is finitely satisfiable, so by the Compactness Theorem it is consistent.
@@
Therefore, extending the action of $\sigma_{\to}$ from polynomials to formulas in the obvious way, as $(\alpha_{jh})_{j\in[1,n],h\in\mathbb{Z}}$ realises $\Theta$, for all $r\leq s\in\mathbb{Z}$ we have, in $\mathcal{L}_{\mathrm{vf}}$,
@@
Recall that, in an $R$-module, a partial positive primitive type (or partial pp-type) is specified by a filter of pp-definable cosets, hence can be written by using at most $\lvert R\rvert+\aleph_{0}$-many formulas (see e.g. [prestModelTheoryModules1988] for model theory of modules). In particular, under our assumptions the $\mathbb{Z}[\sigma]$-module $\mathcal{O}^{\times}$ realises all consistent partial pp-types. By ˜2.14 the short exact sequence of $\mathbb{Z}[\sigma]$-modules $1\to\mathcal{O}^{\times}\to K^{\times}\to\Gamma\to 0$ is pure, hence the existence of a cross-section follows from [prestModelTheoryModules1988, Theorem 2.8]. As standard, a cross-section $s$ induces an angular component $\operatorname{ac}_{s}(x)\coloneqq\operatorname{res}(x/s(v(x)))$, and clearly if $s$ commutes with $\sigma$ then so does $\operatorname{ac}_{s}$, since both $\operatorname{res}$ and $v$ do. The “in particular” part follows from ˜2.2. ∎
@@ 2.2 Good reducts
Let $T_{0}$ be an inductive theory that is not model complete (e.g. the theory of oags with an automorphism, or the theory of fields), and let $T_{1}$ be its Morleyisation (see the references given in ˜1.2). Observe that every model of $T_{0}$ expands to a model of $T_{1}$, and that every model of $T_{1}$ is existentially closed by quantifier elimination. In particular, if $M_{0}\vDash T_{0}$ is not existentially closed, then its natural expansion to $M_{1}\vDash T_{1}$ is existentially closed but has a non-existentially closed reduct.
@@
Let $T$ be the $\mathbb{Z}[\sigma]$-submodule of $L^{\times}$ generated by $K^{\times}$ and $\mathcal{O}(L)^{\times}$, and let $\theta\colon T\rightarrow k(L)^{\times}$ be the unique morphism which coincides with $\operatorname{ac}$ on $K^{\times}$ and with $\operatorname{res}$ on $\mathcal{O}(L)^{\times}$. Enumerate $T$ in a tuple $\bar{b}$. If we prove that the pp-type of $\bar{b}$ in the $\mathbb{Z}[\sigma]$-module $L^{\times}$ is contained in the pp-type of $\theta(\bar{b})$ in $k(L)^{\times}$, then we may conclude by using existential saturation of $k(L)^{\times}$ as a $\mathbb{Z}[\sigma]$-module to apply Point (v) in [prestModelTheoryModules1988, Theorem 2.8], obtaining a morphism $\operatorname{ac}_{L}\colon L^{\times}\rightarrow k(L)^{\times}$ which extends $\theta$.
@@
For all $r\leq s\in\mathbb{Z}$, there is a group homomorphism $\psi_{r,s}\colon\mathbb{G}_{\mathrm{m}}^{n(s+\ell-r+1)}\rightarrow\mathbb{G}_{\mathrm{m}}^{d(s-r+1)}$ satisfying $\psi_{r,s}(\operatorname{orb}_{r,s+\ell}(\alpha))=\operatorname{orb}_{r,s}(a)$, namely we can take the homomorphism which maps $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ to
@@
Let $e=(e_{1},\ldots,e_{d})$ denote the tuple $\xi_{0,0}(\operatorname{orb}_{0,\ell}(\alpha))$. Consider the set $\Omega$ of $\mathcal{L}_{\mathrm{field}}$-formulas over $k(L)$ in variables $(y_{jh})_{j\in[1,n],h\in\mathbb{Z}}$ which says that each finite subtuple $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ is a $k(L)$-generic point of the algebraic variety $V_{r,s}$ defined by the equations $\xi_{r,s}((z_{jh})_{j\in[1,n],h\in[r,s+\ell]})=\theta(\operatorname{orb}_{r,s}(e))$. Note that the genericity requirement makes sense because, as $V_{r,s}(k(L))$ is a coset of $\operatorname{in}_{0}(\ker\xi_{r,s})$ (which does not depend on the choice of an angular component by Lemma˜1.14), it is irreducible by Lemma˜2.11.
To see that $\Omega$ is consistent, consider a finite fragment mentioning formulas in the finite subtuple of variables $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$. As $k(L)^{\times}$ is a divisible abelian group, there is a (possibly non-equivariant) angular component $\operatorname{ac}^{\prime}\colon L\to k(L)$ extending $\theta$. Since, for all $r\leq s$, we have $\operatorname{orb}_{r,s}(e)\in\operatorname{im}(\xi_{r,s})$, and $\operatorname{ac}^{\prime}$ is a multiplicative homomorphism, $V_{r,s}(k(L))$ is not empty.
Hence the set $\Omega$ is realised in some field extension of $k(L)$; let $(\beta_{jh})_{j\in[1,n],h\in\mathbb{Z}}$ be a realisation, and consider the field $k(L)((\beta_{jh})_{j\in[1,n],h\in\mathbb{Z}})$. We now have that, for all $r\leq s$, the finite tuple $(\beta_{j(h+1)})_{j\in[1,n],h\in[r,s+\ell]}$ is generic in the algebraic variety defined by the equations
@@
As $\mathsf{ACVF}_{\mathrm{ac}}$ eliminates quantifiers [^8] This is folklore, and follows from e.g. the much more general [haleviEliminatingFieldQuantifiers2019, Theorem 1]. For a direct proof, use [hilsImaginariesSeparablyClosed2018, Fact 2.4] in combination with [touchardBurdenHenselianValued2023, Observation 1.51]., arguing as in the proof of Proposition˜2.3(1) we see that $K$ is algebraically closed, and in particular closed under taking roots. Suppose that $(L_{0},v_{L_{0}},\sigma_{L_{0}})\supseteq(K,v,\sigma)$ is such that $L_{0}\vDash\exists x\;\varphi(x)$, with $\varphi(x)$ a quantifier-free $\mathcal{L}_{\sigma}(K)$-formula. Let $(L,v_{L},\sigma_{L})$ be an existentially $\aleph_{1}$-saturated existentially closed extension of $L_{0}$. After applying Proposition˜2.18, $L$ extends $K$ as $\mathrm{ac}$-valued difference fields, and $L\vDash\exists x\;\varphi(x)$. Since $\mathcal{L}_{\mathrm{ac},\sigma}\supseteq\mathcal{L}_{\sigma}$ and $(K,v,\sigma,\operatorname{ac})$ is existentially closed, we have $K\vDash\exists x\;\varphi(x)$. ∎
@@
It is well known that the theory $\mathsf{ACVF}_{s}$ of algebraically closed valued fields with a cross-section eliminates quantifiers; see e.g. [Kesting, Theorem 2.2] for a proof for $\mathsf{ACVF}_{s,\iota}$ that can be easily adapted.
@@
Consider the $\mathbb{Z}[\sigma]$-submodule $N$ of $L$ consisting of elements of $L$ with angular component 1; note that the image of $s$ is contained in $N$. Let $\bar{b}$ be a tuple enumerating $\Gamma(K)$. If we prove that the pp-type of $\bar{b}$ in $\Gamma(L)$ is contained in the pp-type of $s(\bar{b})$ in $N$, then we may apply [prestModelTheoryModules1988, Theorem 2.8(v)] and obtain the desired extension of $s$ to $\Gamma(L)$. In other words, we need to show that if $a$ is a finite subtuple of $\bar{b}$, and $A\in\operatorname{Mat}_{d\times n}(\mathbb{Z}[\sigma])$ is such that $A\cdot z=a$ has a solution in $\Gamma(L)$, then $A\cdot z=s(a)$ has a solution in $N$. In fact, it suffices to show that $A\cdot z=a$ has a solution in $\Gamma(K)$, as then if $\gamma$ is such a solution we get that $A\cdot s(\gamma)=s(a)$ and $s(\gamma)\in N$.
This follows from Lemma˜2.20, concluding the proof. ∎
@@
Let $K$ be an existentially closed $\operatorname{ac}$-valued (resp. $s$-valued) difference field. Then $\operatorname{VF}(K)$ is algebraically closed and $(\Gamma,\sigma)$ and $(k,\sigma)$ are existentially closed in their respective categories. In particular, $(K,v,\operatorname{ac})$ (resp. $(K,v,s)$) and $(K,v)$ are existentially closed.
@@
By Proposition˜2.21 it suffices to deal with the $\operatorname{ac}$-valued case. By Corollary˜2.19 $(K,v,\sigma)$ is existentially closed, so we may apply Proposition˜2.3. Alternatively, we may adapt its proof by using quantifier elimination in $\mathsf{ACVF}_{\mathrm{ac}}$. The “in particular” part follows from the above and Proposition˜2.3(1). ∎
@@ 2.3 Bad reducts
Another natural reduct of $(K,v,\sigma)$ to consider is the difference field $(K,\sigma)$. In this subsection we will see that, in contrast to what we saw in Subsection˜2.2, forgetting the valuation does not preserve existential closedness.
Recall that existentially closed difference fields are axiomatisable by a theory denoted by $\mathsf{ACFA}$, in the language $\{+,0,-,\cdot,1,(-)^{-1},\sigma,\sigma^{-1}\}$. The axioms of this theory say that $(K,\sigma)\vDash\mathsf{ACFA}$ iff $K$ is algebraically closed and, for all $n$, all irreducible algebraic subvarieties $V,W$ of $K^{n}$ and $K^{2n}$ respectively such that $W\subseteq V\times\sigma(V)$ and the projections of $W$ to $V$ and to $\sigma(V)$ are dominant, there is $x\in V(K)$ such that $(x,\sigma(x))\in W$. See [Ma, Theorem 7], [CHr, Theorem on p. 3007].
We will see in ˜3.30 that if $k\vDash\mathsf{ACFA}$ has characteristic zero and $\Gamma$ is an existentially closed difference oag, then the Hahn field $k((t^{\Gamma}))$, made into a valued difference field as in ˜1.6, is existentially closed. The following argument, due to Philip Dittmann, shows that it is not existentially closed as a difference field, that is, it is not a model of $\mathsf{ACFA}$.
@@
It is well known (see e.g. [CHr, Proposition 1.2]) that the fixed field of every model of $\mathsf{ACFA}$ is pseudofinite, and in particular not separably closed. By [friedFieldArithmetic2023, Corollary 12.5.5] [^9] In the third edition of the same book (2008), this is Corollary 11.5.5. every pseudo-algebraically closed (in particular, every pseudofinite) nontrivially valued henselian field is separably closed. Since there is a fixed point of nonzero valuation, the induced valuation on the fixed field $K_{0}$ of $K$ is nontrivial. Hence, to obtain a contradiction, it suffices to show that $K_{0}$ is henselian.
We check henselianity directly. Let $f\in\mathcal{O}(K_{0})[x]$ and $a\in\mathcal{O}(K_{0})$ be such that $\operatorname{res}(a)$ is a simple root of $\operatorname{res}(f)$. By henselianity of $K$ there is $b\in\mathcal{O}(K)$ with $f(b)=0$ and v(ba)>0v(b-a)>0. If $\sigma(b)=b$ we are done, so assume this is not the case. As $f$ is over the fixed field, $\sigma(b)$ is a root of $f$. Therefore $(x-b)(x-\sigma(b))$ divides $f$, and since $\operatorname{res}(\sigma(b))=\sigma(\operatorname{res}(b))=\sigma(\operatorname{res}(a))=\operatorname{res}(a)$ this contradicts that $\operatorname{res}(a)$ is a simple root of $\operatorname{res}(f)$. ∎
@@
For notational simplicity, we will only deal with the case $n=2$. Define $P(x)\coloneqq(x-a)(x-\sigma(a))$, observe that it is over the fixed field $K_{0}$, and set $Q(x,y)\coloneqq P(x)P(y)-\varepsilon$. By the Eisenstein criterion (in the version of [friedFieldArithmetic2023, Lemma 2.3.10(b)]) applied to $(K[x])[y]$, the vanishing variety $V$ of $Q$ is absolutely irreducible.
@@
Let us show that every existentially closed valued difference field satisfies the assumptions of Proposition˜2.26. The existence of $\varepsilon$ follows from existential closedness of $\Gamma$ (Proposition˜2.3(2)). To find some $a$ with $v(a)=0$, $\sigma^{2}(a)=a$, and $\operatorname{res}(\sigma(a))\neq\operatorname{res}(a)$, it suffices to take two Gauss extensions (see [EP, Corollary 2.2.2]), let $\sigma$ swap the added points, and apply existential closedness. ∎
@@
The proof above may be adapted, for instance, to the existentially closed isometric case studied in [AvD], but not to the existentially closed strongly contractive case, where the valued field sort is indeed a model of $\mathsf{ACFA}$ [CH, Fact 2.13].